OpenStudy (anonymous):

Three apples cost as much as 4 pears. Three pears cost as much as 2 oranges. How many apples cost as much as 72 oranges ?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

81 apple.... here is what you do:

OpenStudy (anonymous):

cost of 1 apple=x cost of 1 pear=y cost of 1 orange=z

OpenStudy (anonymous):

3 apples = 4 pears (cost ): 3x=4y 3pears = 2 oranges : 3y=2z

OpenStudy (anonymous):

now: one pear y =3/4 x (from the first equation substitute in second: 3y=9/4 x =2z you need 72 oranges - multiply both side by 36: 72z=81x Good?

OpenStudy (anonymous):

I still don't understand the last part everything else is making sence just the last part is confusing.